### v1

# but actually an Hopf, algebra

Abstract
In this paper I would like to show how the Deformed Special Relativity family of models - developed to approach spacetime quantization - can actually be applied to the description of classical cosmology. I use the bicrossproduct basis of $\kappa$-Poincaré algebra to describe photon propagation in deSitter classical General Relativity. I show the Hopf algebraic aspects of deSitter model, and give an explicit physical interpretation of $\kappa$-Poincaré Hopf algebraic features in spacetime. Such an approach allows to unravel some not yet known General Relativistic relations of deSitter cosmology, and reinterpret many future of Quantum Gravity phenomenology as classical properties of maximally symmetric spacetime models.
I Introduction
First of all, this paper is not about Quantum Gravity. This paper is about applying a well defined mathematical model, $\kappa$-Poincaré Hopf algebra [1, 2, 3, 4], very well known and studied in the Quantum Gravity literature, to the study of the signal-exchange between observers in Cosmology and related relativistic features.
$\kappa$-Poincaré is a deformation of Poincaré algebra in which momentum-space has a deSitter-like geometry [5, 6]. This mathematical model can be considered as the backbone of the Deformed Special Relativity (DSR) approach to Quantum Gravity [6, 5, 7, 8, 9, 10]. More or less a decade ago, with my former research group, we showed [11] how one of the most studied phenomenological features of DSR, the relativistic signals energy-dependence (see for instance [12, 13, 14, 15, 16]), was nothing but the dual effect in momentum-space than deSitter cosmological redshift. We showed how just exchanging the cosmological constant $H_{\Lambda}$ with the DSR deformation parameter $\ell$, energy with time, and momenta with particles’ position, the mathematical description of the two effects is basically the same. This interpretation provided us a deeper understanding of DSR approach to Quantum Gravity, since in this framework, the foundation of DSR, the highest achievable energy scale (often identified with Planck energy $E_{P}=\sqrt{\hslash\cdot c^{5}/G}$), becomes in turn nothing but the momentum-space event horizon.
At this point one may ask: Does it mean that $\kappa$-Poincaré is just deSitter mathematical model in momentum-space? The most common and simplistic answer to this question usually is that the resemblance between the two models is just apparent since deSitter one is a Lie algebra, while $\kappa$-Poincaré is in the matter of facts a more complex Hopf algebra [17, 18].
Are we sure about that? I am not questioning $\kappa$-Poincaré being an Hopf algebra, of course it is, but are we so sure that beneath deSitter model it is not possible to unravel any Hopf algebraic aspect? The aim of this paper is precisely to show how also Hopf algebra features like coproduct, antipode and counity can find a compelling interpretation in the framework of curved spacetime relativity. Such an analysis on the one hand is useful to better interpret phenomenology tricky QG phenomenology effects - such as particles vertices in DSR - by comparison with with more familiar phenomena in deSitter; on the other hand Hopf algebraic features could provide a wide set of tools for the study of cosmology and General Relativity.
With respect to our (already cited) previous analysis [11], an important further element to be introduced here is the duality between particles in $\kappa$-Poincaré and observers in deSitter.
This paper is composed by an introductive part: sections II, III; a more theoretical exploration: sections IV, V, VI and VII; and a more phenomenological technical application of DSR features in cosmology in section VIII. In the first two parts I will adopt a set of coordinates in which $c=\hslash=1$, only in Sec. VIII I will switch back to the International System of Units, in order to obtain worldlines in terms of Mpc/year.
In section II a short review of deSitter model is sketched, introducing a few useful elements for the following analysis. In section III are instead briefly reviewed all the fundamental aspects of $\kappa$-Poincaré Hopf algebra, as well as the basic result obtained in our previous analysis [11], also called the lateshift effect (time delay as redshift in momentum-space). In section IV is analyzed the deformed composition law due to non trivial coproduct. Section V and VI explore a not-well-known (and pretty complex) relativistic feature such as backreaction, at first in $\kappa$-Poincaré and then in deSitter. Section VII summarizes the coalgebric aspects of deSitter spacetime and goes into detail on the parallelism between particles in DSR and observers in deSitter. In conclusion in section VIII is addressed the problem of how to apply those theoretical speculations to the study of signal propagation in a phenomenological framework in which astrophisical data are only characterized by measured distance and redshift.
II deSitter spacetime
In order to fix the notation, we resume here the FLRW equations for a homogeneous and isotropic universe can be obtained solving Einstein equation
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R-\Lambda g_{\mu\nu}=\chi T_{\mu\nu}\,,$ (1)
where $\Lambda$ is the cosmological constant, looking for a general solution for the metric, formalized by the line element
$ds^{2}=(dx^{0})^{2}-a(x^{0})^{2}\left(\frac{dr^{2}}{1-kr^{2}}+r^{2}d\Omega^{2}\right)\,,$ (2)
where $a(x^{0})$ is the universe radius of curvature (or scale factor), $k$ is the Gaussian curvature which can be (-1,0,1) for negative, zero or positive curvature and, in order to satisfy the homogeneity and isotropy requirements, one can choose a perfect-fluid like stress-energy tensor:
$T_{\mu\nu}=\left(\begin{array}[]{cccc}\rho&0&0&0\\ 0&-\frac{P}{1-kr^{2}}a^{2}&0&0\\ 0&0&-Pa^{2}r^{2}&0\\ 0&0&0&-Pa^{2}r^{2}\sin^{2}{\theta}\end{array}\right)\,,$ (3)
in which, $\chi=8\pi G/c^{4}$ and $P,\rho$ are the fluid’s energy density and pressure. Einstein equation then reduces to the two FLRW equtions:
$\left(\frac{\dot{a}}{a}\right)^{2}+\frac{k}{a^{2}}=\frac{\chi}{3}\rho+\frac{\Lambda}{3}\,,$ (4)
$2\frac{\ddot{a}}{a}+\frac{k}{a^{2}}+\frac{\dot{a}^{2}}{a^{2}}=\Lambda-\chi P\,.$ (5)
Equations (4) and (5) can easily be coupled, obtaining
$\frac{\ddot{a}}{a}=\frac{\Lambda}{3}-\frac{\chi}{6}(\rho+3P)\,,$ (6)
which in turn, coupled with the first law of thermodynamics for an adiabatic expansion of the universe gives:
$\dot{\rho}=-3\frac{\dot{a}}{a}\left(\rho+P\right)\,.$ (7)
The flat spacetime solution ($k=0$) in absence of matter ($\rho,P=0$) for the above equations implies an exponential scale factor
$a(x^{0})=e^{\pm\sqrt{\frac{\Lambda}{3}}x^{0}}\,,$ (8)
which specializes the family of Friedmann spacetimes to the maximally symmetric de Sitter spacetime (see for an interesting insight [19]) with $H_{\Lambda}=\sqrt{\Lambda/3}$.
DeSitter spacetime is a maximally-symmetric deformation of Minkowski spacetime, in which the familiar Lorentz group in four dimensions $L(c)=SO(1,3)$ will be associated with the de Sitter group $dS(c,H_{\Lambda})$, which is the Lorentz group $SO(1,4)$ in five dimensions. Therefore a maximally-symmetric four-dimensional spacetime manifold can be constructed by identifying it both metrically and geometrically with the quotient space $dS(c,H_{\Lambda})/L(c)$. This procedure is equivalent to the identification of this spacetime manifold with the de Sitter one $dS(1,3)$ which is the one sheeted four-dimensional hyperboloid of equation
$\eta_{AB}X^{A}X^{B}=(X^{0})^{2}-(X^{1})^{2}-(X^{2})^{2}-(X^{3})^{2}-(X^{4})^{2}=-\frac{1}{H_{\Lambda}^{2}}\,,$ (9)
where $\eta_{AB}$ is the five-dimensional Minkowskian metric with signature (+,,,,)(+,-,-,-,-)( + , - , - , - , - ). Due to the presence of curvature, in general, no privileged class of reference frames exists on the Minkowskian hyperboloid and any choice of a local coordinate system may be in principle acceptable for the description of physical phenomena; however in order to recognize at first sight a solution with exponential scale factor (8), we will focus on the flat set of coordinates:
$X^{0}$ $=$ $\frac{1}{H_{\Lambda}}\sinh(H_{\Lambda}x^{0})+\frac{H_{\Lambda}}{2}e^{H_{\Lambda}x^{0}}|x|^{2}$
$X^{i}$ $=$ $x^{i}e^{H_{\Lambda}x^{0}}$ (10)
$X^{4}$ $=$ $\frac{1}{H_{\Lambda}}\cosh(H_{\Lambda}x^{0})-\frac{H_{\Lambda}}{2}e^{H_{\Lambda}x^{0}}|x|^{2}\,,$
Since the line-element should be invariant under any set of coordinates, one can find the relation between the five-dimensional $\eta$ and the four-dimensional $g$ metric imposing
[…]
and therefore
$g_{\mu\nu}(x)=\eta_{AB}\frac{\partial X^{A}}{\partial x^{\mu}}\frac{\partial X^{B}}{\partial x^{\nu}}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&-e^{2H_{\Lambda}x^{0}}&0&0\\ 0&0&-e^{2H_{\Lambda}x^{0}}&0\\ 0&0&0&-e^{2H_{\Lambda}x^{0}}\\ \end{array}\right)\,.$ (12)
In order to find the spacetime symmetry generators $\Pi_{0}$, $\Pi_{i}$, ${\cal N}_{i}$ and ${\cal R}_{i}$ in one can use the Killing differential equation
$\partial_{\alpha}g_{\mu\nu}\Xi^{\alpha}+g_{\alpha\nu}\partial_{\mu}\Xi^{\alpha}+g_{\mu\alpha}\partial_{\nu}\Xi^{\alpha}=0\,.$ (13)
Such an equation, using the metric (12) reduces to the differential equations system:
$\left\{\begin{array}[]{l}\partial_{0}\Xi^{0}=0\\ \partial_{i}\Xi^{0}=e^{2H_{\Lambda}x^{0}}\partial_{0}\Xi^{i}\\ \partial_{i}\Xi^{i}=-H_{\Lambda}\Xi^{0}\\ \partial_{i}\Xi^{j}+\partial_{j}\Xi^{i}=0\end{array}\right.\;\;\;,\;\;\;$ (14)
whose solutions are:
$\Xi^{0}(x)$ $=$ $a^{0}+\delta_{jk}\xi^{j}x^{k}\,,$ (15)
$\Xi^{i}(x)$ $=$ $-a^{0}H_{\Lambda}x^{i}+a^{i}-\epsilon_{ijk}\alpha^{j}x^{k}+\xi^{i}\left(\frac{1-e^{-2H_{\Lambda}x^{0}}}{2H_{\Lambda}}+\frac{H_{\Lambda}}{2}|x|^{2}\right)-H_{\Lambda}x^{i}\delta_{jk}\xi^{j}x^{k}\,,$ (16)
where $a^{\mu},\alpha^{i},\xi^{i}$ are the spacetime transformations parameters.Once found the Killing vectors explicit expression, using the total charge definition $Q\equiv\Xi^{\mu}(x)p_{\mu}$, where the $p_{\mu}$ are the locally-flat quadrimomenta, it is possible to identify the symmetry generators representations
$\Pi_{0}=p_{0}-H_{\Lambda}x^{i}p_{i}\;,\;\;\;\Pi_{i}=p_{i}\,,\mathcal{R}_{i}=\epsilon_{ijk}x^{j}p_{k}$ (17)
$\mathcal{N}_{i}=x^{i}p_{0}+\left(\frac{1-e^{-2H_{\Lambda}x^{0}}}{2H_{\Lambda}}+\frac{H_{\Lambda}}{2}|x|^{2}\right)p_{i}-H_{\Lambda}x^{i}x^{k}p_{k}\,,$ (18)
and finally the deSitter Lie algebra:
$[\Pi_{0},\Pi_{i}]=-iH_{\Lambda}\Pi_{i}\;,\;\;\;[\Pi_{i},\Pi_{j}]=0\;,$ (19)
$[\mathcal{R}_{i},\Pi_{0}]=0\;,\;\;\;[\mathcal{R}_{i},\Pi_{j}]=i\epsilon_{ijk}\Pi_{k}\;,\;\;\;[\mathcal{R}_{i},\mathcal{N}_{j}]=i\epsilon_{ijk}\mathcal{N}_{k}\;,\;\;\;[\mathcal{R}_{i},\mathcal{R}_{j}]=i\epsilon_{ijk}\mathcal{R}_{k}$ (20)
$[\mathcal{N}_{i},\Pi_{0}]=-i\left(\Pi_{i}-H_{\Lambda}{\cal N}_{i}\right)\;,\;\;\;[\mathcal{N}_{i},\Pi_{j}]=-i\Pi_{0}\delta_{ij}-iH_{\Lambda}\epsilon_{ijk}\mathcal{R}_{k}\;,\;\;\;[\mathcal{N}_{i},\mathcal{N}_{j}]=-i\epsilon_{ijk}\mathcal{R}_{k}\;.$ (21)
Whose firts Casimir is
${\cal H}=\left(\Pi_{0}+x^{i}\Pi_{i}\right)^{2}-e^{-2H_{\Lambda}x^{0}}|\Pi|^{2}=p_{0}^{2}-e^{-2H_{\Lambda}x^{0}}|p|^{2}\,.$ (22)
It is also possible at this point to characterize spacetime coordinates transformations under boost, through the differential equations arising from the relation
$\frac{dx^{\mu}}{d\xi^{i}}=-i[\mathcal{N}_{i},x^{\mu}]\,.$ (23)
With the right choice of coordinates, imposing to be aligned with the point to be boosted, id est imposing for instance $i=1$ and the point original coordinates to be $(\bar{x}^{0},\bar{x}^{1},0,0)$, after some cumbersome calculations, one can find:
$\left\{\begin{array}[]{l}x^{0}(\xi)=\frac{1}{H_{\Lambda}}\ln\left(e^{H_{\Lambda}\bar{x}^{0}}\left((\cosh\frac{\xi}{2}-H_{\Lambda}\bar{x}^{1}\sinh\frac{\xi}{2})^{2}-e^{-2H_{\Lambda}\bar{x}^{0}}\sinh^{2}\frac{\xi}{2}\right)\right)\\ x^{1}(\xi)=\frac{e^{2\xi}-1-e^{2H_{\Lambda}\bar{x}^{0}}e^{-\xi}(1-H_{\Lambda}\bar{x}^{1})^{2}+e^{2H_{\Lambda}\bar{x}^{0}}(1+H_{\Lambda}\bar{x}^{1})^{2}}{H_{\Lambda}\left(e^{2H_{\Lambda}\bar{x}^{0}}(1+H_{\Lambda}\bar{x}^{1}+e^{\xi}(1-H_{\Lambda}\bar{x}^{1})^{2})-(e^{\xi}-1)^{2}\right)}\end{array}\right.\,.$ (24)
The same procedure for quadrimomenta gives
$\left\{\begin{array}[]{l}p_{0}(\xi)=-\bar{p}_{0}\cosh\left(\ln\left(\frac{1-e^{\xi}-e^{H_{\Lambda}\bar{x}^{0}}\left(1+H_{\Lambda}\bar{x}^{1}+e^{\xi}(1-H_{\Lambda}\bar{x}^{1})\right)}{1-e^{\xi}+e^{H_{\Lambda}\bar{x}^{0}}\left(1+H_{\Lambda}\bar{x}^{1}+e^{\xi}(1-H_{\Lambda}\bar{x}^{1})\right)}\right)\right)+e^{-H_{\Lambda}\bar{x}^{0}}\bar{p}_{1}\sinh\left(\ln\left(\frac{1-e^{\xi}-e^{H_{\Lambda}\bar{x}^{0}}\left(1+H_{\Lambda}\bar{x}^{1}+e^{\xi}(1-H_{\Lambda}\bar{x}^{1})\right)}{1-e^{\xi}+e^{H_{\Lambda}\bar{x}^{0}}\left(1+H_{\Lambda}\bar{x}^{1}+e^{\xi}(1-H_{\Lambda}\bar{x}^{1})\right)}\right)\right)\\ p_{1}(\xi)=\frac{e^{-\xi}e^{-2H_{\Lambda}\bar{x}^{0}}}{4}\left((e^{\xi}-1)^{2}\bar{p}_{1}-e^{2H_{\Lambda}\bar{x}^{0}}\left(1+H_{\Lambda}\bar{x}^{1}+e^{\xi}(1-H_{\Lambda}\bar{x}^{1})\right)\left(2(e^{\xi}-1)\bar{p}_{0}-\bar{p}_{1}\left(1+H_{\Lambda}\bar{x}^{1}+e^{\xi}(1-H_{\Lambda}\bar{x}^{1})\right)\right)\right)\end{array}\right.\,.$ (25)
Both (24) and (25) in the limit $H_{\Lambda}\rightarrow 0$ reduce to the classical Lorentz transformations
$\left\{\begin{array}[]{l}x^{0}(\xi)=\bar{x}^{0}\cosh(\xi)-\bar{x}^{1}\sinh(\xi)\\ x^{1}(\xi)=-\bar{x}^{0}\sinh(\xi)+\bar{x}^{1}\cosh(\xi)\end{array}\right.\;,\;\;\;\left\{\begin{array}[]{l}p_{0}(\xi)=\bar{p}_{0}\cosh(\xi)-\bar{p}_{1}\sinh(\xi)\\ p_{1}(\xi)=-\bar{p}_{0}\sinh(\xi)+\bar{p}_{1}\cosh(\xi)\end{array}\right.$ (26)
It is pretty easy to verify that both the line element obtained from (11):
$\Delta s^{2}=\frac{4}{H_{\Lambda}^{2}}\sinh\left(\frac{H_{\Lambda}}{2}x^{0}\right)^{2}-e^{H_{\Lambda}x^{0}}|x|^{2}\,,$ (27)
and the first Casimir ${\cal H}$ introduced in (22), are invariant under the $H_{\Lambda}$-deformed Lorentz transformations (24) and (25).
Long story short, deSitter group is a General Relativity induced deformation to Poincaré group. Other kinds of Poincaré algebra deformations have been applied to the study of spacetime quantization. Such models are known as Doubly Special Relativity (DSR) models [7, 8, 9, 10, 13], they are mostly based on the so called $\kappa$-Poincaré Hopf algebra.
III Hopf Algebras for dummies
The most significant difference between Hopf algebras and the more familiar Lie algebras, is the presence of a non-trivial coproduct, which is an element of the coalgebric sector [17, 18]. While an algebra is a vectorial space on a field $F$, in which we have defined the concept of:
$\begin{array}[]{ll}Product&{\cal M}:{\cal A}\otimes{\cal A}\rightarrow{\cal A}\\ Unity&\eta:F\rightarrow{\cal A}\end{array}$ (28)
a coalgebra is a vectorial space ${\cal A}$ on a field $F$, provided of the linear maps:
$\begin{array}[]{ll}Coproduct&\Delta:{\cal A}\rightarrow{\cal A}\otimes{\cal A}\\ Counity&\epsilon:{\cal A}\rightarrow F\end{array}$ (29)
The coalgebras theory is essentially dual to the one of the algebras. In fact, if ${\cal A}$ is a coalgebra, then is dual space ${\cal A}*$ (the space of the linear functionals $\Phi:{\cal A}\rightarrow F$) inherits a punctual algebra structure induced by $\Delta$ and $\epsilon$.
An Hopf algebra is a vectorial space ${\cal A}$ on a field $F$ which is simoultaneusly both an algebra and a coalgebra, and which is provided by a further linear map called antipode:
${\cal S}:{\cal A}\rightarrow{\cal A}$ (30)
whose role is analogous to the inverse element in the link between an algebra and a coalgebra. A Lie algebra can be generalised into an Hopf algebra by mean of its universal enveloping algebra $U(g)$, trivially defining coproduct, counity and antipode:
$\begin{array}[]{l}\Delta(X)=X\otimes\mathbb{1}+\mathbb{1}\otimes X\\ \epsilon(X)=0\\ {\cal S}(X)=-X\end{array}$ (31)
Hopf algebras are a very powerful tool for the study of Planck-scale physics, in fact, interpreting the Planck length $L_{P}=\sqrt{\hbar G/c^{3}}$ as an observer-independent scale rather than a coupling constant, might bring to a direct conflict with the basic principles of Special Relativity, in which energies and lengths cannot be observer-independent. To overcome this issue one can rely to a theory with two observer-independent scales, the already mentioned DSR models.
III.1 $\kappa$-Poincaré
The Deformed Special Relativity approach to Quantum Gravity assumes the possibility to formalize the spacetime quantization through some sort of coordinates noncommutativity, such as
$[\chi^{\mu},\chi^{\nu}]=\Theta^{\mu\nu}+\Xi_{\gamma}^{\mu\nu}\chi^{\gamma}\,.$ (32)
The $\Theta^{\mu\nu}\neq 0$ case is the so called ”canonical” approach to coordinates noncommutativity, while specializing on the case $\Xi_{\gamma}^{\mu\nu}\sim\ell$ one obtains the $\kappa$-Minkowski coordinates commutation rules:
$[\chi^{0},\chi^{i}]=-i\ell\chi^{i}\,,$ (33)
where $\ell$ is a parameter with the dimensions of a length that we expect to be of the order of the inverse of Planck scale $\ell\sim 1/M_{P}$.
The request for the relation (33) to be invariant under a 10-generators spacetime symmetry group, imposes to take into account an $\ell$-deformed algebra. Given the parallelism between (33) and the first relation of (19), the most studied deformed algebras taken into account are deSitter and ant-deSitter inspired Hopf algebras (see for instance [1, 2, 3, 4, 20])).
The natural candidate for this theory’s algebra of symmetries is a family of quantum (Hopf) algebras generalizing the standard Poincaré algebra. In particular the DSR proposal can be realized in the framework of quantum $\kappa$-Poincaré algebra [5, 6]. This algebra contains a natural deformation parameter of mass $\kappa\sim M_{P}$ by construction (for our purposes we will prefer to consider as deformation parameter a lenght scale $\ell$, defined as $\ell=1/\kappa$).
$\kappa$-Poincaré Hopf algebra bicrossproduct basis is characterized [21] by the following commutation relations:
$[p_{i},p_{j}]=0\,,\;\;\;[{\cal R}_{i},{\cal R}_{j}]=-i\epsilon_{ijk}{\cal R}_{k}\,,\;\;\;[{\cal R}_{i},p_{j}]=-i\epsilon_{ijk}p_{k}\,,$
$\left[{\cal N}_{i},p_{0}\right]=-ip_{i}\,,\;\;\;[{\cal N}_{i},p_{j}]=-i\delta_{ij}\left(\frac{1-e^{-2\ell p_{0}}}{2\ell}+\frac{\ell}{2}|p|^{2}\right)+i\ell p_{i}p_{j}\,,$ (34)
$[{\cal N}_{i},{\cal N}_{j}]=i\epsilon_{ijk}{\cal R}_{k}\,,\;\;\;[{\cal R}_{i},{\cal N}_{j}]=-i\epsilon_{ijk}{\cal N}_{k}\,,$
and the following coproducts and antipodes
$\Delta p_{0}=p_{0}\otimes\mathbb{1}+\mathbb{1}\otimes p_{0}\,,$ (35)
$\Delta p_{i}=p_{i}\otimes\mathbb{1}+e^{-\ell p_{0}}\otimes p_{i}\,,$ (36)
${\cal S}(p_{0})=-p_{0}\,,\;\;\;{\cal S}(p_{i})=-e^{\ell p_{0}}p_{i}\,,$ (37)
$\Delta{\cal N}_{i}={\cal N}_{i}\otimes\mathbb{1}+e^{-\ell p_{0}}\otimes{\cal N}_{i}-i\ell\epsilon_{ijk}p_{j}\otimes{\cal R}_{k}\,,$ (38)
${\cal S}({\cal N}_{i})=-e^{\ell p_{0}}{\cal N}_{i}+i\ell\epsilon_{ijk}p_{j}{\cal R}_{k}\,,$ (39)
where the ${\cal R}_{i}$ are the undeformed rotations and ${\cal N}_{i}$ is the boost generator which can be represented as
${\cal N}_{i}=\chi^{0}p_{i}+\chi^{i}\left(\frac{1-e^{-2\ell p_{0}}}{2\ell}+\frac{\ell}{2}|p|^{2}\right)=x^{0}p_{i}+x^{i}\left(\frac{1-e^{-2\ell p_{0}}}{2\ell}+\frac{\ell}{2}|p|^{2}\right)-\ell x^{j}p_{j}p_{i}\,.$ (40)
The boost can be both represented in terms of noncommuattive $\chi^{\alpha}$ or commutative coordinates $x^{\alpha}$, the map linking the two coordinatizations
$\chi^{0}=x^{0}-\ell x^{i}p_{i}\,,\;\;\;\chi^{j}=x^{j}\,,$ (41)
is well known in the literature (see exempli gratia [16, 22]) and often used to work with an undeformed symplectic sector $[p_{\alpha},x^{\beta}]=-i\delta_{\alpha}^{\beta}$. The duality between noncommutative coordinates and the deSitter physical momenta $\Pi_{\mu}$ has been explored in [11], finding that the map (41) has the same form as the one relating local momenta $p_{\mu}$ with the conserved charges (17) that one can find solving the deSitter Killing equation.
The interpretation of the physical consequences of this kind of deformed algebra is still debated, in some cases they seem to imply features like particles in-vacuo dispersion (id est momentum-dependent photon velocity) [7, 8, 13, 15, 16] and deformed non-linear composition laws for momenta, strictly connected with the coproduct nontriviality.
The relative locality interpretation [23, 24] portrays those features as momentum-space curvature effects: the momentum-dependent massless particles velocity has been succesfully described [11] as some sort of dual redshift effect or lateshift, while the modified composition law is in general identified at first orders as
$(p\oplus q)_{\alpha}\equiv p_{\alpha}+q_{\alpha}-\ell Z_{\alpha}^{\beta\gamma}p_{\beta}q_{\gamma}+\dots$ (42)
where the physical interpretation of $Z$ have not been univocally identified and is still debated [^1] It have been often postulated the identification between $Z$ and the momentum-space connections $Z_{\alpha}^{\beta\gamma}\equiv\Gamma_{\alpha}^{\beta\gamma}$, however such an assumption imposes to take into account some kind of momentum-space torsion [23, 25], which we won’t take into account here, since deSitter spacetime has no torsion..
Moreover another feature characterizes the modified composition law, namely the backreaction [26, 25, 27, 28]. This is a property of the composition law under boost: if we have a particle with four-momentum $k_{\alpha}$ decaying into two particles with momenta $p_{\alpha}$ and $q_{\alpha}$, one would like that the composition law observed in a boosted reference frame with boost parameter $\xi$ to have the same form as the one at rest, however:
$k^{\prime}_{\alpha}(\xi)\neq(p^{\prime}(\xi)\oplus q^{\prime}(\xi))_{\alpha}\,.$ (43)
The reason for this weird behaviour is that in those models momenta ”backreact” on the boost parameter itself $\xi\vartriangleleft(p,q)$ [29], according to the boost nontrivial coproduct (38). This last aspect of $\kappa$-Poincaré algebra may be a little tricky to explain, however, since the lateshift approach is so effective in describing Hopf algebra features just as momentum-space curvature effects with a deSitter-like metric, with no need for torsion or nonmetricity, why not to try a similar strategy to formalize the modified composition law and the backreaction?
As long as one can find a duality relation between time delay effects and redshift, it may be possible also to find a similar liaison between interacting particles in an $\ell$ deformed momentum-space and observers in deSitter spacetime. The identification of a satisfying physical interpretation of such features using deSitter spacetime as inspiration is the heart of this article and will be extensively addressed in sec. IV. Before elaborating any further about the Hopf-algebraic aspects of deSitter geometry we may need to clarify a little bit more how $\kappa$-Poincaré momentum-space and deSitter spacetime are related by reviewing some well established DSR features and their physical interpretation.
III.2 Parallelism between $\kappa$-Poincaré phenomenology and deSitter Redshift
$\kappa$-Poincaré curved momentum-space algebra belongs to a well populated family of quantum groups of spacetime symmetries [30]. In this case we will focus on the deSitter-like geometry, starting from [31, 29, 32] the Hamiltonian obtained by integrating the momentum-space invariant line-element
[…]
where the $\zeta$ is the momentum-space metric and where the momentum-space geodesics $P(s)$ are defined by the geodesic equation:
$\ddot{P}_{\alpha}+\Gamma_{\alpha}^{\beta\gamma}\dot{P}_{\beta}\dot{P}_{\gamma}=0\,.$ (45)
The deformed Hamiltonian (44) is related to the deSitter-like momentum-space metric
$\zeta^{\alpha\beta}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&-e^{-2\ell P_{0}}&0&0\\ 0&0&-e^{-2\ell P_{0}}&0\\ 0&0&0&-e^{-2\ell P_{0}}\end{array}\right)\,,$ (46)
through the on-shell relation ${\cal H}=m^{2}$ where the mass is the invariant momentum-space line-element defined as
[…]
where the $P(s)$ are momentum-space geodesics which satisfy $P_{\alpha}(0)=0\,,P_{\alpha}(1)=p_{\alpha}$ and whose explicit expressions
$P_{0}(s)$ $=$ $-\frac{1}{\ell}\ln(1+(e^{\ell p_{0}}-1)s)\,,$ (48)
$P_{i}(s)$ $=$ $\frac{e^{\ell p_{0}}p_{i}s}{1+(e^{\ell p_{0}}-1)s}\,,$ (49)
can be obtained solving the geodesic equations (45), in which the connection $\Gamma$ is the usual Christoffel symbol in momentum-space
$\Gamma_{\lambda}^{\mu\nu}=\frac{1}{2}\zeta_{\lambda\sigma}\left(\frac{\partial}{\partial p_{\mu}}\zeta^{\sigma\nu}+\frac{\partial}{\partial p_{\nu}}\zeta^{\sigma\mu}-\frac{\partial}{\partial p_{\sigma}}\zeta^{\mu\nu}\right)\,,$ (50)
with no torsion or nonmetricity:
$\Gamma_{0}^{00}=\Gamma_{0}^{01}=\Gamma_{0}^{10}=\Gamma_{i}^{00}=\Gamma_{i}^{ii}=0\,$ (51)
$\Gamma_{0}^{ii}=-\ell e^{-2\ell p_{0}}\;,\;\;\;\Gamma_{1}^{10}=\Gamma_{i}^{0i}=-\ell\,.$ (52)
We can then write down the Hamiltonian as the algebra first Casimir:
${\cal H}=\frac{4}{\ell^{2}}\sinh^{2}\left(\frac{\ell}{2}p_{0}\right)-e^{\ell p_{0}}|p|^{2}\,,$ (53)
this leads to a modified Hamiltonian framework, characterized by a rich phenomenology of deformation effects for a free massless particle propagating in spacetime, formalized by the (classical) Hamilton equations:
$\dot{x}^{\beta}=\{{\cal H},x^{\beta}\}\,.$ (54)
Using this relation it easy to find the explicit worldlines expression:
[…]
If we imagine to have two massless particle emitted at the same time at an observer’s (Alice) origin $(x^{0}_{A}=0,x^{1}_{A}=0)$ with different energies: an high energetic $p_{0}^{h}$ and a low energetic one $p_{0}^{l}$, the worldlines (55) imply that a translated observer (Bob) will not receive those two particles at the same time. In fact in Bob’s frame
$x^{\alpha}_{B}\simeq x^{\alpha}_{A}-a^{\beta}\{p_{\beta},x^{\alpha}\}\,,$ (56)
where $(a^{0},a^{1})$ are the translation coefficients, the two massless particles will arrive at the spacial origin $x^{1}=0$ at two different times, whose difference characterizes the time delay
$\Delta x_{B}^{0}=a^{0}(e^{\ell(p_{0}^{h}-p_{0}^{l})}-1)\simeq\ell a^{0}(p_{0}^{h}-p_{0}^{l})\,.$ (57)
Relative Locality [23, 24, 15, 16] is a classical framework in which both quantum and spacetime curvature effects are turned off, i.e. $\hslash\rightarrow 0$ and $G\rightarrow 0$. At the same time however Planck-scale deformations are still present as modification to the classical Hamiltonian mechanics since in this limit $\ell\sim M_{P}=\sqrt{G/\hbar}\neq 0$. In this limit then all the aforementioned Hopf-algebraic features are treated as remnants of a quantum regime in a classical framework with deformed Hamiltonian dynamics with a curved momentum-space. For this reason all the features so far formalized using commutation relations can be also expressed as nontrivial Poisson brackets, following the well known relation $\left[A,B\right]=i\hslash\{A,B\}$ and the coordinate set taken into account will be the locally flat commutative one $x^{\alpha}$ instead of the noncommuative one $\chi^{\beta}$.
III.3 Lateshift
The duality between the $\kappa$-Poincaré time-delay and the spacetime redshift effect in deSitter framework is pretty easy to point out [11]. First of all one can notice that given (53) and (55), in the massless case we have
[…]
The Relativity of locality of $\kappa$-Minkowski noncommutative spacetime model can be defined observing that in such a framework events that are coincident and occur in the origin of a given observer Alice are objectively coincident for all observers that share Alice’s origin (all observers which are purely boosted with respect to Alice), but may not appear to be coincident to an observer Bob far from Alice. Locality is still objective, but the abstraction of ”distant locality” is lost.
According to (55), two photons with energy $\bar{p_{0}}$ and $\tilde{p}_{0}$ emitted at Alice origin $(\bar{x}^{0}_{A}=0,\bar{x}^{1}_{A}=0)$ will propagate along the worldlines
$x^{1}_{A}=e^{\ell\bar{p}_{0}}x^{0}_{A}\,,$ (59)
$x^{1}_{A}=e^{\ell\tilde{p}_{0}}x^{0}_{A}\,.$ (60)
Given the coordinate transformations (56) between Alice and Bob
$\bar{x}^{0}_{B}=\bar{x}^{0}_{A}-a^{0}\,,$ (61)
$\bar{x}^{1}_{B}=\bar{x}^{1}_{A}-a^{0}\,,$ (62)
and the relation between the translation parameters $a^{1}=e^{\ell\bar{p}_{0}}a^{0}$ (which can be found defining Bob’s origin as the intersection between Bob and the first worldline), Bob will express the two worldlines as
$x^{1}_{B}=e^{\ell\bar{p}_{0}}(x^{0}_{B}-a^{0})-e^{\ell\bar{p}_{0}}a^{0}\,,$ (63)
$x^{1}_{B}=e^{\ell\tilde{p}_{0}}(x^{0}_{B}-a^{0})-e^{\ell\bar{p}_{0}}a^{0}\,.$ (64)
Therefore the time delay between the arrival af the two particles at Bob’s spatial origin ($x^{1}_{B}=0$, where the detector lies) can be expressed as
$\frac{\Delta x^{0}_{B}}{\bar{x}^{0}_{B}}=1-e^{-\ell\Delta p_{0}}\,,$ (65)
where $\Delta p_{0}=\tilde{p}_{0}-\bar{p}_{0}$.
The duality between this effect and the redshift in momentum-space can be highlighted by asking what happends in this framework in an exact specular situation: if in $\kappa$-Poincaré we are interested in the time delay arising between two particles with different energies emitted at the same time, in the deSitter case we will take into account two photons emitted at different times with the same energy. Let’s consider then a first photon emitted with a generic energy $p_{0}^{A}$ from some standard candle (Alice), being observed by a translated observer Bob at a distance $(a^{0},a^{1})$. The energy of the photon in Bob’s frame will be determined by the well known redshift formula in deSitter:
$\bar{p}_{0}^{B}=e^{-H_{\Lambda}a^{0}}p_{0}^{A}\,.$ (66)
After some time $\Delta x^{0}$, in an expanding spacetime the distance between the two observers will of course increase and the energy $\tilde{p}_{0}^{B}$ measured by Bob will be further redshifter by
$\tilde{p}_{0}^{B}=e^{-H_{\Lambda}(a^{0}+\Delta x^{0})}p_{0}^{A}=e^{-H_{\Lambda}\Delta x^{0}}\bar{p}_{0}^{B}\,.$ (67)
Therefore the redshift effect can be characterized as
$\frac{\Delta p_{0}^{B}}{\bar{p}_{0}^{B}}=1-e^{-H_{\Lambda}\Delta x^{0}}\,.$ (68)
Quod erat demonstrandum. It is worth to mention that in order to find this duality between the $\kappa$-Poincaré momentum-space and deSitter spacetime we didn’t have to assume any torsion or nonmetricity in the geometry of our manifolds, on the other hand looking for an expression for the deformed momentum composition law in terms of parallel-transported vectors may lead to more complex scenarios that we may not need in a lateshift-like approach.
IV Nontrivial composition law for momenta and deSitter reference frames
While the deformed Hamiltonian (53) formalizes the $\kappa$-Poincaré modified dynamics of particles, the coproducts such as (36) and (35) give us some indications about the particles’ interacition. In fact the coproduct can also be expressed [33, 28] as
$\Delta p_{\mu}={\cal D}_{\mu}(p\otimes\mathbb{1},\mathbb{1}\otimes p)\,,$ (69)
where the function ${\cal D}_{\mu}(p,q)$ describes a deformed composition of momenta ${\cal D}_{\mu}(p,q)=(p\oplus q)_{\mu}$.
In other words, since translations in $\kappa$-Minkowski spacetime, defined from the $\kappa$-Minkowski commutation rules between coordinates (33) can be represented [2, 1] through simple derivatives:
$p_{0}\equiv-i\frac{\partial}{\partial\chi^{0}}\;,\;\;\;p_{0}\equiv-i\frac{\partial}{\partial\chi^{0}}\,,$ (70)
we can obtain the total momentum and energy of two interacting particles, calculating the the eigenvalues of those operators acting on two (time-to-the-right) plane waves $\Phi(\chi)=e^{ip_{i}\chi^{i}}e^{ip_{0}\chi^{0}}$ and $\Psi(\chi)=e^{iq_{i}\chi^{i}}e^{iq_{0}\chi^{0}}$
Since our coordinates are noncommutative we should make use of the Baker-Campbell-Hausdorff formula which implies that in case $[X,Y]=sY$ then $e^{X}e^{Y}=e^{e^{s}Y}e^{X}$. Moving all the time coordinates to the right we have that one can define the composition rule for the particles’ momenta as
$e^{ip_{i}\chi^{i}}e^{ip_{0}\chi^{0}}e^{ik_{i}\chi^{i}}e^{ik_{0}\chi^{0}}=e^{i(p\oplus k)_{i}\chi^{i}}e^{i(p\oplus k)_{0}\chi^{0}}\,.$ (71)
Therefore since we have that
$[p_{0}\chi^{0},k_{i}\chi^{i}]=[p_{0},k_{i}\chi^{i}]\chi^{0}+p_{0}[\chi^{0},k_{i}\chi^{i}]=p_{0}k_{i}[\chi^{0},\chi^{i}]=p_{0}k_{i}(-i\ell\chi^{i})\,,$ (72)
using the BKH-formula we can identify the composition rules of this model as
$k_{\mu}=\left\{\begin{array}[]{l}(p\oplus q)_{0}=p_{0}+q_{0}\\ (p\oplus q)_{i}=p_{i}+e^{-\ell p_{0}}q_{i}\end{array}\right.\,,$ (73)
which is in agreement with the definition (69).
This kind of emergy and momenta composition rules modifications have in general striking implications for the macroscopical limit which are currently not entirely comprehended. The discussion on how to compose those effect through some sort of Planck-scale-sensitive statistical mechanics have lead to the so called soccerball problem [34, 35] which tries to address the issue of machrosopical manifestations which of course are not detected in table-top experiments. There is some literature about how to solve the soccerball problem using curved momentum-space formalism [36]. In this paper we would like to give a few more insight about such an interpretation, showing how an explicit duality can be identified between modified composition laws for momenta in $\kappa$-Poincaré and deSitter spacetime features.
IV.1 Composition law in deSitter spacetime
As well as it is possible to define a strict duality between redshift and the time-delay effect, it is also possible to explain $\kappa$-Poincaré non trivial composition law of momenta as a composition of vectors in deSitter spacetime, however since the boost acts on the composition law nonlocally, id est one needs to take into account the backreaction effect to ensure the invariance of interaction vertices under boost transformations, we won’t assume this composition law to be the sum of parallel transported vectors as in subsection 42. We can imagine the two vectors to lie at the origins of two translated observers Alice and Bob, which are connected by a photon worldline
$x_{A}^{1}(x_{A}^{0})=\frac{1-e^{-H_{\Lambda}x^{0}_{A}}}{H_{\Lambda}}$ (74)
so that both their origins sit in different point of such worldline. Let’s now say that Alice wants to know what are the coordinates of some vector $\vec{k}$ that in terms of Bob coordinates is $\vec{q}_{@B}=(b^{0},b^{1})$, knowing that their origins are connected by some other vector that Alice expresses as $\vec{p}_{@A}=(a^{0},a^{1})$ (see Figure 1).
Figure 1: Alice tries to express the components of vector $\vec{k_{@A}}$ in terms of vector $\vec{p}_{@A}$ which lies in her reference frame and $\vec{q}_{@B}$ whose components are defined by the translated observer Bob.
How can Alice formalize the vector $\vec{k}_{@A}$ as the composition of the coordinates of $\vec{p}_{@A}$ and $\vec{q}_{@B}$? For the class of observers reached in their spacetime origin by the signal emitted in Alice’s spacetime origin $(x^{1}_{B}(x^{0}_{B}=0)=0)$, it is possible to define the following coordinate transformations under translation:
$x^{0}_{B}$ $=$ $\bar{x}^{\alpha}\Pi_{\alpha}\vartriangleright x^{0}_{A}=x^{0}_{A}-a^{0}$ (75)
$x^{1}_{B}$ $=$ $\bar{x}^{\alpha}\Pi_{\alpha}\vartriangleright x^{1}_{A}=e^{H_{\Lambda}a^{0}}\left(x^{1}_{A}-\frac{1-e^{-H_{\Lambda}a^{0}}}{H_{\Lambda}}\right)\,,$ (76)
where the action of the symmetry generators $\Pi_{0}=p_{1}-H_{\Lambda}x^{1}p_{1}$, $\Pi_{1}=p_{1}$ on the phase space functions is represented through the ordinary action-to-the-right of Lie groups
$\bar{a}{\mathcal{G}}\vartriangleright A=\sum_{n=0}^{\infty}\frac{1}{n!}\{\bar{a}{\mathcal{G}},\{\ldots\{\bar{a}{\mathcal{G}},A\}\ldots\}\,,$ (77)
in which the $\bar{a}$ are the coordinate transformation parameters.
It is then easy to find how Alice will have to transform $\vec{q}_{@B}$ coordinates provided by Bob:
$k^{0}_{@A}$ $=$ $a^{0}+b^{0}$ (78)
$k^{1}_{@A}$ $=$ $e^{-H_{\Lambda}a^{0}}b^{1}+\frac{1-e^{-H_{\Lambda}a^{0}}}{H_{\Lambda}}\,.$ (79)
Moreover, since Bob’s origin sits on the worldline that originates in Alice’s one, we can use (74) to find how Alice will express vector $\vec{k}_{@A}$ in terms of $\vec{p}_{@A}$ and $\vec{q}_{@B}$ components:
$\vec{k}_{@A}=\vec{p}_{@A}\oplus\vec{q}_{@B}=(a^{0},a^{1})\oplus(b^{0},b^{1})=\vec{p}_{@A}+\vec{q}_{@A}=\left(a^{0}+b^{0}\,,\,a^{1}+b^{1}e^{-H_{\Lambda}a^{0}}\right)\,.$ (80)
We can observe that (80) reproduces in spacetime the composition rules (73) expressed for $\kappa$-Poincaré momentum space. We don’t have in this case an explicit expression in terms of connections as in [29, 23] but a much simpler formalism that only depends on the metric of spacetime. We don’t need to explicitly parallel transport any vector but we express $\vec{k}_{@A}$ directly as the sum of vectors $\vec{p}_{@A}$ in Alice’s frame and $\vec{q}_{@B}$ in Bob’s one.
Therefore while on the one hand there’s a well established duality between worldlines in deSitter and Casimir in $\kappa$-Poincaré, as well as between redshift and DSR time-delay effect, we have now found that, on the other hand, different particles in $\kappa$-Poincaré behave like different reference frames in deSitter spacetime.
V Backreaction in $\kappa$ Poincaré
Let’s now on $\kappa$-Poincaré peculiar relativistic features. In relativistic particle physics it is, of course, of paramount importance to coherently formalize how different observers will describe some particles interaction and which are the relations between the observables that they will respectively measure. For instance in the simple case described in (73), one would expect that since $k_{\alpha}=(p\oplus q)_{\alpha}$, then a boosted observer would have $k_{\alpha}^{\prime}=(p^{\prime}\oplus q^{\prime})_{\alpha}$, where we have denoted with $k^{\prime},p^{\prime},q^{\prime}$ the momenta measured by the latter. However in the $\kappa$-Poincaré framework this relativistic prescription is not as intuitive, since:
$k^{\prime}_{\alpha}(\xi)\neq(p^{\prime}(\xi)\oplus q^{\prime}(\xi))_{\alpha}\,.$ (81)
In fact, letting the boost in the direction 1 act as
$A^{\prime}=\mathcal{N}_{(1)}\vartriangleright A=A+\xi_{1}\{\mathcal{N}_{(1)},A\}+\frac{\xi_{1}^{2}}{2!}\{\mathcal{N}_{(1)},\{\mathcal{N}_{(1)},A\}\}+\frac{\xi_{1}^{3}}{3!}\{\mathcal{N}_{(1)},\{\mathcal{N}_{(1)},\{\mathcal{N}_{(1)},A\}\}\}+...\,,$ (82)
it is not difficult to verify that the $\kappa$ Poincaré boost acts differently on the right and on the left term of relation (81). Since in the following section we will be interested mostly by the non-ultrarelativistic regime of the deSitter transformations (24), for sake of simplicity we can here show this different behaviour at first order in the rapidity parameter $\xi_{1}$. Therefore, while the boost of the momenta composition is
$k^{\prime}_{1}$ $=$ $k_{1}+\xi_{1}\left(\frac{1-e^{-2\ell k_{0}}}{2\ell}-\ell k_{1}^{2}+\frac{\ell}{2}|k|^{2}\right)=$ (83)
$=$ $(p\oplus q)_{1}^{\prime}=p_{1}+e^{-\ell p_{0}}q_{1}+\xi_{1}\left(\frac{1-e^{-2\ell(p_{0}+q_{0})}}{2\ell}-\ell(p_{1}+e^{-\ell p_{0}}q_{1})^{2}+\frac{\ell}{2}|p+e^{-\ell p_{0}}q|^{2}\right)\,,$
on the other hand the composition of the boosted momenta gives
$(p^{\prime}\oplus q^{\prime})_{1}$ $=$ $p_{1}^{\prime}+e^{-\ell p_{0}^{\prime}}q_{1}^{\prime}=$
$=$ $p_{1}+e^{-\ell p_{0}}q_{1}+\xi_{1}\left(\frac{1-e^{-2\ell p_{0}}+e^{-\ell p_{0}}-e^{-\ell(p_{0}+2q_{0})}}{2\ell}-\ell(p_{1}^{2}+e^{-\ell p_{0}}p_{1}q_{1}+e^{-\ell p_{0}}q_{1}^{2})+\frac{\ell}{2}(|p|^{2}+e^{-\ell p_{0}}|q|^{2})\right)\,.$
Confronting those two results we can verify relation (81) just calculating
$(p\oplus q)_{1}^{\prime}-(p^{\prime}\oplus q^{\prime})_{1}=\xi_{1}\left(\frac{e^{-2\ell(p_{0}+q_{0})}}{2\ell}(e^{\ell p_{0}}-1-e^{\ell(p_{0}+2q_{0})}(1+\ell^{2}(2q_{1}^{2}-|q|^{2}+2(p_{2}q_{2}+p_{3}q_{3})))+e^{2\ell q_{0}}(1-\ell^{2}(2q_{1}^{2}+|q|^{2}))\right)\,.$ (85)
At first sight this feature may suggest some kind of unreliability of the $\kappa$-Poincaré relativistic framework, since apparently two boosted observers could not have any way to come to an agreement about the result of some particles’ interaction. However this is just a misleading interpretation based on the naive assumption that the generators should act trivially on plane waves as expected from a formalization based on simple lie algebras (31):
$\mathcal{N}_{i}\vartriangleright(e^{ip_{\mu}x^{\mu}}e^{iq_{\nu}x^{\nu}})=(\mathcal{N}_{i}\vartriangleright e^{ip_{\mu}x^{\mu}})e^{iq_{\nu}x^{\nu}}+e^{ip_{\mu}x^{\mu}}(\mathcal{N}_{i}\vartriangleright e^{iq_{\nu}x^{\nu}})\,,$ (86)
while, according to the nontrivial boost coproduct (38), the boost action on coupled plane-waves should be
$\mathcal{N}_{i}\vartriangleright(e^{ip_{\mu}x^{\mu}}e^{iq_{\nu}x^{\nu}})=(\mathcal{N}_{i}\vartriangleright e^{ip_{\mu}x^{\mu}})e^{iq_{\nu}x^{\nu}}+(e^{-\ell P_{0}}\vartriangleright e^{ip_{\mu}x^{\mu}})(\mathcal{N}_{i}\vartriangleright e^{iq_{\nu}x^{\nu}})-i\ell\epsilon_{ijk}(P_{j}\vartriangleright e^{ip_{\mu}x^{\mu}})({\cal R}_{k}\vartriangleright e^{iq_{\nu}x^{\nu}})\,.$ (87)
The action at (87) should seem not easily manageable, however in [25, 29] (see also [37] for a clearer explanation) was suggested to re-express such feature as some sort of backreaction of momenta on the boost parameter $\xi\vartriangleleft(p,q)$, as already formalized in [26]. The vertex conservation in $\kappa$-Poincaré is in general satisfied by
$k^{\prime}_{\alpha}(\xi)=(p^{\prime}(\xi\vartriangleleft q)\oplus q^{\prime}(\xi\vartriangleleft p))_{\alpha}\,.$ (88)
In the bicrossproduct basis, however, at first order in the rapidity parameter $\xi$ equation (88) reduces to:
$k^{\prime}_{\alpha}(\xi)\simeq(p^{\prime}(\xi)\oplus q^{\prime}(\xi e^{-\ell p_{0}}))_{\alpha}\,.$ (89)
If one was ineterested to study such feature in the ultrarelativistic regime, a generalization of this last equation working well beyond the first order is provided in [25, 26], which is:
$\xi\vartriangleleft p_{i}=2\arcsin\left(\frac{e^{-\ell p_{0}}\sinh\frac{\xi}{2}}{\sqrt{(\cosh\frac{\xi}{2}-\ell p_{i}\sinh\frac{\xi}{2})^{2}-(e^{-\ell p_{0}}\sinh\frac{\xi}{2})^{2}}}\right)\,.$ (90)
However this formalization gives a coherent description of particles’ vertices under boost in $\kappa$-Poincaré, the problem related to the dependence of the reaction result to particles’ ordering, already present in the non-boosted vertex, becomes even more problematic: if $p\oplus q\neq q\oplus p$, and therefore $p^{\prime}(\xi)\oplus q^{\prime}(\xi\vartriangleleft p)\neq q^{\prime}(\xi)\oplus p(\xi\vartriangleleft q)$, when measuring $k$ and $k^{\prime}$ should different observers agree on which particle comes first in the momenta composition? How should nature know about the particles momenta ordering that we are writing on paper? In the quantum-gravity framework this problem have been addressed on many different angles [38, 31, 39, 40], and still is an open issue, on the other hand in a classical curved-spacetime framework such ordering mechanisms should not imply the same kind of issues, since there are clear rules governing the transformations between different observers in General Relativity.
VI Backreaction in deSitter
In Section IV.1 we have defined three vectors $\vec{p},\vec{q},\vec{k}$ as expressed in Alice’s (@$A$) and Bob’s (@$B$) coordinate sets:
$\begin{array}[]{c}p_{@A}=(a^{0},a^{1})\;\;\;\;\;,\;\;\;\;\;q_{@B}=(b^{0},b^{1})\,,\\ q_{@A}=(b^{0},b^{1}e^{-H_{\Lambda}a^{0}})\;\;\;\;\;,\;\;\;\;\;k_{@A}=(a^{0}+b^{0},a^{1}+b^{1}e^{-H_{\Lambda}a^{0}})\,.\end{array}$ (91)
We also defined a composition law $\oplus$ so that those three vectors are forced to close a vector sum along a curved de Sitter worldline (see Figure 1):
$\vec{k}_{@A}=\vec{p}_{@A}\oplus\vec{q}_{@B}\,.$ (92)
We will define the observer Camilla as the boosted observer along the direction $1$, sharing it’s origin with Alice. The vectors’ components will transform under boost according to (24). Since a boost is well defined only in the origin we can directly boost only the vectors whose origin is in Alice’s and Camilla’s origin
$p_{@C}=(p_{@A})^{\prime}(\xi)\;,\;k_{@C}=(k_{@A})^{\prime}(\xi)$ (93)
Boosting directly Bob’s coordinates won’t allow-us to close the vector sum in Camilla’s frame, as clearly represented by the Magenta arrow in Figure 2.
Figure 2: The vector sum $\vec{k}=\vec{p}\oplus\vec{q}$ as expressed in the frame at rest (in black) and in the boosted one (in red). The Magenta vector represents $q^{\prime}$ without taking into account the backreaction on the rapidity parameter $\xi$. The values fixed for the parameters are $H=1$, $a^{0}=b^{0}=1$ and $\xi=0.5$.
Therefore for what concerns $\vec{q}$, whose origin lies in Bob, we will have to take into account also an effect on the rapidity, then
$q_{@D}=(q_{@B})^{\prime}(\xi\vartriangleleft p_{@A})\,,$ (94)
where given our choice of coordinates the action $\xi\vartriangleleft p_{@A}$ will be defined as the spacetime dual of equation (90). Vector $q_{@D}$ won’t be the expression of $q_{@B}$ in Camilla’s system, but in a translated one (say Diego). therefore if we want to compose $p$ and $q$ we need to remind that in curved spacetime we will have
$(k_{@A})^{\prime}(\xi)=k_{@C}=p_{@C}+q_{@C}=p_{@C}\oplus q_{@D}=(p_{@A})^{\prime}(\xi)\oplus(q_{@B})^{\prime}(\xi\vartriangleleft p_{@A})\,.$ (95)
Let’s now try to perform this calculation one step at the time. For sake of simplicity we will limit to first order in $\xi$, which in a cosmological framework (non-ultrarelativistic regime) is perfectly justified. We will also assume the boost to be alined with the signal direction ($x^{2},x^{3}=0$). Therefore:
$(k_{@A})^{\prime}(\xi)\simeq k_{@A}-i\xi[\mathcal{N}_{1},k_{@A}]=\left\{\begin{array}[]{l}a^{0}+b^{0}-\xi(a^{1}+b^{1}e^{-H_{\Lambda}a^{0}})\\ a^{1}+b^{1}e^{-H_{\Lambda}a^{0}}+\xi\left(\frac{e^{-2H_{\Lambda}(a^{0}+b^{0})}-1}{2H_{\Lambda}}+\frac{H_{\Lambda}}{2}(a^{1}+b^{1}e^{-H_{\Lambda}a^{0}})^{2}\right)\end{array}\right.\,,$ (96)
$(p_{@A})^{\prime}(\xi)\simeq p_{@A}-i\xi[\mathcal{N}_{1},p_{@A}]=\left\{\begin{array}[]{l}a^{0}-\xi a^{1}\\ a^{1}-\xi\left(\frac{1-e^{-2H_{\Lambda}a^{0}}}{2H_{\Lambda}}-\frac{H_{\Lambda}}{2}(a^{1})^{2}\right)\end{array}\right.\,,$ (97)
$(q_{@B})^{\prime}(\xi\vartriangleleft p_{@A})\simeq q_{@B}-i\xi e^{-H_{\Lambda}a^{0}}[\mathcal{N}_{1},p_{@A}]=\left\{\begin{array}[]{l}b^{0}-\xi e^{-H_{\Lambda}a^{0}}b^{1}\\ b^{1}-\xi e^{-H_{\Lambda}a^{0}}\left(\frac{1-e^{-2H_{\Lambda}b^{0}}}{2H_{\Lambda}}-\frac{H_{\Lambda}}{2}(b^{1})^{2}\right)\end{array}\right.\,.$ (98)
Gathering all together, and taking into account that we lie along a deSitter worldline $x^{1}=(1-e^{-H_{\Lambda}x^{0}})/H_{\Lambda}$, it is straightforward to verify equation (95).
Then, while in the $\kappa$-Poincaré approach to quantum gravity a clear physical reason for the ordering dependence of momenta vertices is still debated, in the deSitter framework the interpretation is pretty easy. The backreaction in curved spacetime formalizes the need to ”close” the vector sum between vectors laying on different observers frames (see Fig.2). The ordering is relative to the observer who wants to analyze and confront signals originating from different translated sources, while taking into account apparent velocity discrepancies.
VII In summarising de Sitter Hopf algebra
It is common knowledge that $\kappa$-Poincaré is a deSitter-inspired curved momentum-space Hopf algebra. What is not-so-common knowledge, however, is that also in deSitter framework we can recognize an Hopf-algebra structure. In fact, in summarizing what we obtained so far, the deSitter deformed Lorentz generators $\mathcal{N}_{i},\mathcal{R}_{j}$ together with coordinates $x^{\mu}$ close an Hopf algebra with algebric sector
$[x^{\mu},x^{\nu}]=0\;,\;\;\;[\mathcal{R}_{i},\mathcal{R}_{j}]=-i\epsilon_{ijl}\delta^{lk}\mathcal{R}_{k}\;,\;\;\;[\mathcal{N}_{i},\mathcal{N}_{j}]=i\epsilon_{ijl}\delta^{lk}\mathcal{R}_{k}\;,$
$[\mathcal{N}_{i},x^{0}]=-i\delta_{ij}x^{j}\;,\;\;\;[\mathcal{N}_{i},x^{j}]=-i\delta_{i}^{j}\left(\frac{1-e^{-2H_{\Lambda}x^{0}}}{2H_{\Lambda}}+\frac{H_{\Lambda}}{2}|x|^{2}\right)+iH_{\Lambda}x^{k}x^{j}\delta_{k}^{i}\;,$ (99)
$[\mathcal{R}_{i},x^{0}]=0\;,\;\;\;[\mathcal{R}_{i},x^{j}]=-i\epsilon_{ilk}\delta^{lj}x^{k}\;,\;\;\;[\mathcal{R}_{i},\mathcal{N}_{j}]=-i\epsilon_{ijl}\delta^{lk}\mathcal{N}_{k}\,,$
and whose coalgebric sector is dual to the $\kappa$-Poincaré one:
$\Delta x^{0}=x^{0}\otimes\mathbb{1}+\mathbb{1}\otimes x^{0}\,,$
$\Delta x^{i}=x^{i}\otimes\mathbb{1}+e^{-H_{\Lambda}x^{0}}\otimes x^{i}\,,$
${\cal S}(x^{0})=-x^{0}\,,\;\;\;{\cal S}(x^{i})=-e^{H_{\Lambda}x^{0}}x^{i}\,,$ (100)
$\Delta{\cal N}_{i}={\cal N}_{i}\otimes\mathbb{1}+e^{-H_{\Lambda}x^{0}}\otimes{\cal N}_{i}-iH_{\Lambda}\epsilon_{mjk}\delta^{mi}x^{j}\otimes{\cal R}_{l}\delta^{lk}\,,$
${\cal S}({\cal N}_{i})=-e^{H_{\Lambda}x^{0}}{\cal N}_{i}+iH_{\Lambda}\epsilon_{mjk}x^{j}\delta^{mi}\delta^{lk}{\cal R}_{k}\,,$
Such an Hopf algebra defines also an algebra between the different observers frames, so that
$\Delta x^{\mu}(x_{@A},x_{@B})=(x_{@A}\oplus x_{@B})^{\mu}\,,$ (101)
which is associative given the coassociativity axiom of Hopf algebras:
$((x_{@A}\oplus x_{@B})\oplus x_{@C})^{\mu}=(x_{@A}\oplus(x_{@B}\oplus x_{@C}))^{\mu}\,.$ (102)
As well as the coproduct defines the composition law, on the other hand the antipode defines the inversion $\ominus$:
${\cal S}(x^{\mu})_{@A}=\ominus x^{\mu}_{@A}\,.$ (103)
The last element we need to define is the origin of the observer’s framework which in turn can be defined through the deSitter Hopf algebra counity :
$x^{\mu}(\underline{0})=\epsilon(x^{\mu})=((\ominus x_{@A})\oplus x_{@A})^{\mu}\,.$ (104)
The spacetime observers’ algebra is then equipped with a neutral object $\underline{0}$, a composition rule $\oplus$ and the inversion $\ominus$, for which the associativity property holds. Composition law and inversion can be intuitively described as sliding along the worldline connecting different observers. Of course the worldlines’ shape depends on the initial choice of coordinates (10) from which in turn we can define a different Killing equation, different symmetry generators and of course a different algebra. A different set of coordinates in the observers’ framework will then imply a change in the basis of the Hopf algebra we will rely on, which in turn will bring to a different definition of the composition law $\oplus$ and its inversion $\ominus$, just as expected, since also the worldlines would change form accordingly. This means that there is a well defined relation between spacetime diffeomorfism and the choice of the Hopf algebra basis.
This idea of sliding along the worldlines helps-us to develop an intuition also for what concerns the boost and the backreaction: while in flat spacetime one can immagine that a boosted observer will perceive all the other observers to be boosted relatively to him with the same parameter $\xi$ (since in special relativity each observer will transform indipendently from the others), in curved spacetime this won’t be true anymore and the boost performed in some observer’s origin will act differently on the others, according to their distance along the worldline, so that if one observer at rest expresses some point position as the composition of a few vectors
$r_{@A}=p_{@A}\oplus q_{@B}\oplus k_{@C}\,,$ (105)
a boosted observer $D$ in the origin of $A$ will have:
$r_{@D}=r_{@A}^{\prime}(\xi)$ $=$ $p_{@A}^{\prime}(\xi)\oplus q_{@B}^{\prime}(\xi\vartriangleleft p_{@A})\oplus k_{@C}^{\prime}(\xi\vartriangleleft p_{@A}\vartriangleleft q_{@B})=$ (106)
$=$ $p_{@A}^{\prime}(\xi)\oplus q_{@B}^{\prime}(\xi\vartriangleleft p_{@A})\oplus k_{@C}^{\prime}(\xi\vartriangleleft(p_{@A}\oplus q_{@B}))\,.$
Of course the identification
$\xi\vartriangleleft p_{@A}\vartriangleleft q_{@B}=\xi\vartriangleleft(p_{@A}\oplus q_{@B})\,,$ (107)
is implied by the associativity of $\oplus$, but it also expresses that the right action on $\xi$ grows the more the boosted observer origin is far away from the origin of a signal along the signal’s worldline. In our context this may imply that an observer will perceive closer objects as more boosted than further ones, which may have interesting implications from the phenomenological point of view.
VIII Nontrivial composition law and cosmological redshift
The application of the previously introduced mathematical framework to the study of cosmological and astrophisical observations, would, of course, require further exploration of more complex non-maximally-symmetric spacetime models. Such investigation might need the development of ad hoc tecniques (also inspired to QG literature see for instance [41]), in order to study the Hopf-algebraic features of signals in a non-empty universe.
Since the aim of this paper is the introduction of the physical interpretation of some basic and general Hopf-algebraic features, it may be better to leave such a complex task to future investigations. However a simple way to introduce a more concrete element in our analysis is the study of redshift coproduct and its practical implications.
In cosmology the sygnals time-of-emission is an inference of the model. The actual observables are, on the other hand, distance (from the receiver point of view) and redshift $z$. The definition of redshift is the very well known $z=(\lambda_{observed}-\lambda_{emitted})/\lambda_{emitted}$, which in our simple model translates into
$z=e^{Hx^{0}}-1\,.$ (108)
Since in Section IV we used spacetime translations to provide a physical interpretation to coproduct, it may make sense now to try a similar approach to obtain a redshift coordinate coproduct. Since using the definition (108) we have:
$b^{\alpha}\{\Pi_{\alpha},z\}=b^{0}He^{Hx^{0}}=b^{0}H(1+z)\,,$ (109)
applying iteratively the procedure at eq. (77) to Alice observer’s redshift coordinate $z_{a}$, we obtain:
$b^{\alpha}\Pi_{\alpha}\vartriangleright z_{a}=z_{a}+z_{b}\,e^{Ha^{0}}\,,$ (110)
where of course $z_{b}=e^{Hb^{0}}-1$ and $a^{0}$ is the inferred Alice time coordinate in the observer framework. Therefore redshift coproduct can be defined as:
$\Delta z=z\otimes\mathbb{1}+e^{H_{\Lambda}x^{0}}\otimes z\,.$ (111)
However since, as we already stated, time-coordinate (which is the only one that scales linearly) can be inferred only after the definition of a model, basing on it an analysis aimed to the study of cosmological signals, may lead to a fundamental inconsistency. Therefore it is useful to redefine coordinate composition (and their related coproducts) in terms of physical observables:
$z_{a}\oplus_{z}z_{b}=e^{Hx_{a}^{0}+Hx_{b}^{0}}-1=z_{a}+z_{b}e^{Hx_{a}^{0}}=z_{a}+z_{b}(1+z_{a})\,,$ (112)
and of course for what concerns space coordinates:
$x_{a}^{i}\oplus_{x}x_{b}^{i}=x_{a}^{i}+x_{b}^{i}e^{-Hx_{a}^{0}}=x_{a}^{i}+\frac{x_{b}^{i}}{1+z_{a}}\,,$ (113)
which lead to:
$\begin{array}[]{c}\Delta z=z\otimes\mathbb{1}+(1+z)\otimes z\,,\\ \Delta x^{i}=x^{i}\otimes\mathbb{1}+(1+z)^{-1}\otimes x^{i}\,.\end{array}\,.$ (114)
As in Section IV, also (112) and (113) express vector coordinates composition along a worldline $x^{i}(z)$, the only difference with the $x^{i}(x^{0})$ case, is that now both coordinates evolve nontrivially. The way to deal with this minor issue is to immagine a series of translated locally flat observers, with flat translation parameters, $(t,x,z)\rightarrow(\epsilon^{0},\epsilon^{1}=c\cdot\epsilon^{0},\epsilon^{z}=(H_{\Lambda}/c)\epsilon^{1})$, and express the point of view of the observer at the origin as the total nontrivial composition of $n$ locally flat frames coordinates. Any point $P(z,x^{i})$ along a photon worldline can be then expressed as:
$P(z,x^{i})=(\epsilon^{z}_{A},\epsilon^{i}_{A})\oplus_{z,x}(\epsilon^{z}_{B},\epsilon^{i}_{B})\oplus_{z,x}\dots\oplus_{z,x}(\epsilon^{z}_{\Omega},\epsilon^{i}_{\Omega})\,,$ (115)
where, of course, the indices $\oplus_{z,x}$ below the composition sign, remind us that it stands for different rules for $z$ and $x^{i}$, according to (112) and (113).
Figure 3: On the left $n$ locally Special relativistic vectors in a z/x graphic are composed according to nontrivial coproducts (114). The larger $n$, the more accurate is the reconstruction of the deSitter worldlines $x^{i}(z)$. The values of $n$ represented are (3,4,10,20,30,1000). On the right a reconstructed deSitter model through nontrivial vector composition, with $(\epsilon^{z}\sim 0.126,\epsilon^{1}\sim 13\text{Mpc})$ and $H_{\Lambda}=3\cdot 10^{-10}\text{years}^{-}1$, is showed juxtaposed to the SN Ia data from [42].
As we can observe in Figure 3, the locality of the flatness requirement is of paramount importance since the composition of not-narrow-enough vectors fails to collapse to the correct deSitter worldline. However no complex numerical simulations are required in order to obtain satisfactory results: a few hundreds elements composition is practically Indistinguishable from the curve $x^{1}(z)$.
The fact that the nontrivial composition of a large number of locally Special Relativistic vectors is able to re-express a photon’s worldline can be also showed in a more formal way. In fact the precondition of locally flatness make possible to write the additional term in (113) as:
$\frac{\delta x^{i}}{1+z}\simeq\frac{c\delta x^{0}}{1+z}\,.$ (116)
Since the $x^{0}$ coordinate scales linearly, the composition (115) of infinite locally flat elements can be expressed as an integral, which allows us to recover the worldline definition in terms of redshift:
[…]
With the definition of the coproduct of $z$ we are now provided with all the elements, in terms of physical interpretation of mathematical symbols, needed to explore cosmological observations using the formalism of $\kappa$-Poincaré; previously employed only to approach the Quantum Gravity problem. Of course a lot of issues related with spacetime symmetries breakdown in presence of energy and matter (bright or dark at will) should be addressed, but, as already stated, this is led to future explorations.
IX Conclusions
In summarizing, we have seen that an Hopf algebraic aspects of deSitter model, although never aknowledged in the literature, can be found in many commonly and not-so-commonly studied cosmological frameworks.
Coproduct, for instance, is very naturally embodied in deSitter model as signal propagation in curved spacetime, and can be seen as a different way to reconstruct the already known photon worldlines shapes under different hypotheses about the configuration of the universe. Other Hopf algebraic features like backreaction are way less obvious to recognise in the general framework of deSitter cosmology, however, once correctly interpreted, may provide a different angle to analyze relativistic properties of curved spacetimes.
If on the one hand coproduct for coordinates may be seen just as a different way to sum or integrate portions of spacetime, the study of the deSitter ”coalgebric sector” may hold a few surprises when studying boost composition or the interaction between boosts and translations. In particular a more in-depth study of boost coproduct and in general the action of deSitter boost on different observers, subject to the relativistic relations summarized in Section VII, may not be insignificant, in a field of study in which reconstructing objects redshift plays a crucial role.
Of course this article is intended just as a previous exploration regarding the application of Quantum Groups $\grave{a}$ la Majid for the study of General Relativity. Such a mathematical framework provides a wide set of tools to effectively dissect relativistic phenomena. If coproduct and backreaction are that interesting, what can we say about Weyl maps, star product or Drinfeld twists? We know what $H_{\Lambda}$ is in cosmologic framework, but what may be its interpretation in a Quantum Algebra framework?
In parallel with the mathematical analysis a further evolution of this work may be an intense phenomenological exploration, extending this approach to non-maximally-symmetric models of the universe (such as $\Lambda CDM$) in order to apply the mathematical model to real, actual cosmological data, studying wether this novel tools are able to suggest new features or new ways to analyze the old ones.
There is still a lot of work to be done in order to understand if this kind of in-depth-study may actually be productive from a physical perspective or not. However what I would like to communicate to all those working on Quantum Gravity and DSR is that there is some digging to do and that you have the right shovels.
Acknowledgements.
I would like to thank Antonino Marcianó for his suggestions and insightful conversations.
I aknowledge support from the Huzhou Rihua Information Technology Co., Ltd and Zhejiang Kelin Talent Development Co., Ltd in whose offices this paper was partially written.
I also thank an unknown (allegedly Russian) hacker for giving me the necessary time and peace of mind to set up this work, by devastating the informatic system of my previous workplace.